OpenStudy (anonymous):

if i know an even plus even equals and even number. then can i suppose for contradiction that an even plus and even equals and odd, and give an example like 2+2 is even thus concluding that and even plus even must be even?

jimthompson5910 (jim_thompson5910):

you can prove it directly any even number is of the form 2*k where k is an integer so let x = 2*m y = 2*n m,n are integers so that means x and y are even numbers

jimthompson5910 (jim_thompson5910):

add them up x+y = 2*m + 2*n = 2*(m+n) = 2*p where p = m+n is an integer (summing two integers yields another integer) so because x+y = 2*p this means x+y is even too

OpenStudy (anonymous):

yes, thank you but is the contradiction statement true, i believe it is not but dont know why

jimthompson5910 (jim_thompson5910):

you could do a proof by contradiction, but it would follow basically the same steps as shown above step 1) assume `even` plus `even` is `odd` step 2) go through the work shown above step 3) step 2 will conclude with x+y being even, which contradicts the assumption in step 1

OpenStudy (anonymous):

okay thank you

jimthompson5910 (jim_thompson5910):

no problem